OpenStudy (juana02):

can someone simply this fraction? √ ̅125 ------ √ ̅5

OpenStudy (danjs):

\[\sqrt{125} / \sqrt{5}\] That?

OpenStudy (juana02):

yes

OpenStudy (danjs):

you can start by seeing if any of those 2 nubers are perfect squares

OpenStudy (alexandervonhumboldt2):

so multlply by sqrt5/sqrt5

OpenStudy (danjs):

multiply the both top and bottom by square root 5

OpenStudy (alexandervonhumboldt2):

(sqrt125*5)/5

OpenStudy (danjs):

\[\frac{ \sqrt{125}\sqrt{5} }{ \sqrt{5}\sqrt{5} }\]

OpenStudy (alexandervonhumboldt2):

you get sqrt625/5

OpenStudy (alexandervonhumboldt2):

=25/5=5

OpenStudy (alexandervonhumboldt2):

5

OpenStudy (danjs):

awesome answer alex! you are so smart. Now to see if the actual question person knows how to do that.

OpenStudy (dtan5457):

lol ^

OpenStudy (alexandervonhumboldt2):

thx then could you medal me? @DanJS

OpenStudy (dtan5457):

@AlexandervonHumboldt2 i learned it the hard way, it's tempting to just give the answer but seriously it's actually more fun to know the person knows what they are doing..

OpenStudy (zzr0ck3r):

\(\frac{\sqrt{125}}{\sqrt{5}}=\frac{5^\frac{3}{2}}{5^\frac{1}{2}}=5^{\frac{3}{2}-\frac{1}{2}}=5^\frac{2}{2}=5\)

OpenStudy (alexandervonhumboldt2):

my father teached me thats when i was 5

OpenStudy (dtan5457):

i think you missed the sarcastic statement by @DanJS

OpenStudy (zzr0ck3r):

or \(\frac{\sqrt{125}}{\sqrt{25}}=\sqrt{\frac{125}{5}}=\sqrt{25}=5\)

OpenStudy (danjs):

Do you recall @juana02 that the square root of a number , is the same as the number to the 1/2 power? \[\sqrt{n} = n ^{1/2}\]

OpenStudy (danjs):

for example \[\sqrt{5} = 5^{1/2}\]

OpenStudy (danjs):

Or are you just doing square roots of perfect square numbers so far?

OpenStudy (danjs):

in the case where, these two expressions are equal, you can make it one square root \[\frac{ \sqrt{125} }{ \sqrt{5} } = \sqrt{\frac{ 125 }{ 5 }}\]

OpenStudy (danjs):

ok i will be on for awhile, let me know when you return @juana02